tarah2024:

Laura rents a movie for a flat fee of $2.00 plus an additional $0.50 for each night she keeps the movie. Choose the cost function that represents this scenario if x equals the number of nights Laura has the movie.

Mercury:

answer choices: c(x) = 2.00x + 0.50 c(x) = 2.00 + 0.50x c(x) = 2.50x c(x) = (2.00 + 0.50)x

Mercury:

total cost: flat fee + (cost per night)(# of nights) fill in the appropriate numbers/variables from the problem.